Given that R = 25 cm, r = 10 cm and that the machine requires an effort of 500 N to lift a 100 kg mass, determine the:

i. velocity ratio

ii. weight of the 100 kg mass

iii. mechanical advantage

iv. efficiency

v. useful work done in lifting the mass through a height of 65 cm.

vi. law of the machine


1
Expert's answer
2021-06-14T15:00:46-0400

Gives

R=25cm

r=10cm

Part (a)

Velocity ratio

"=\frac{E_d}{L_d}=\frac{25}{10}=2.5"

Part (b)

Waight of 100kg mass

"w=mg=100\times9.8=980N"

Part(c)

Machenical advantages

"M.A=\frac{Load}{effort}"

"M.A=\frac{980}{500}=1.96"

Part(d)

Efficiency "=\frac{M.A}{V.R}=\frac{1.96}{2.5}=0.784"

Part(e)

Work done by lifting mass(m=100kg)

h=0.65m

W=mgh

"W=100\times9.8\times0.65=637J"

Part(f)

Law of machine gives idea about the load and effort relation of a machine. It also tells us

about the friction present in the machine under no load condition.

Law of machine can be used to find the load when effort is known and effort when load is known

and also maximum mechanical advantage.